### v1

# Pairing mechanism and superconductivity in pressurized La5Ni3O11

Abstract
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen in pressurized La3Ni2O7 has sparked a surge of exploration of high-$T_{c}$ superconductors in the Ruddlesden-Popper (RP) phase nickelates. More recently, the RP phase nicklate La5Ni3O11, which hosts layered structure with alternating bilayer and single-layer NiO2 planes, is reported to accommodate SC under pressure, exhibiting a dome-shaped pressure dependence with highest $T_{c}\approx 64$ K, capturing a lot of interests. Here, using density functional theory (DFT) and random phase approximation (RPA) calculations, we systematically study the electronic properties and superconducting mechanism of this material. Our DFT calculations yield a band structure including two nearly decoupled sets of sub- band structures, with one set originating from the bilayer subsystem and the other from the single-layer one. RPA-based analysis demonstrates that SC in this material occurs primarily within the bilayer subsystem exhibiting an $s^{\pm}$ wave pairing symmetry similar to that observed in pressurized La3Ni2O7, while the single-layer subsystem mainly serves as a bridge facilitating the inter-bilayer phase coherence through the interlayer Josephson coupling (IJC). Since the IJC thus attained is extremely weak, it experiences a prominent enhancement under pressure, leading to the increase of the bulk $T_{c}$ with pressure initially. When the pressure is high enough, the $T_{c}$ gradually decreases due to the reduced density of states on the $\gamma$-pocket. In this way, the dome-shaped pressure dependence of $T_{c}$ observed experimentally is naturally understood.
The discovery of superconductivity (SC) with critical temperature $T_{c}\approx 80$ K in the Ruddlesden-Popper (RP) bilayer nickelate La3Ni2O7 under pressure bilayernature has sparked significant experimental and theoretical interest in the field bilayermodel ; huomodulation ; zhang2023electronic ; lechermann2023electronic ; luo2024high ; wu2024superexchange ; cpl_41_7_077402 ; shilenko2023correlated ; yang2023interlayer ; zhang2023trends ; christiansson2023correlated ; shen2023effective ; oh2023type ; liu2023s ; liao2023electron ; yang2023possible ; PhysRevB.110.235155 ; kaneko2024pair ; ouyang2024absence ; heier2024competing ; zhang2024structural ; zhang2024electronic ; tian2024correlation ; ryee2024quenched ; zhang2024strong ; ni_spin_2025 ; lu2024interlayer ; qu2024bilayer ; yang2024strong ; fan2024superconductivity ; sakakibara2024possible ; cao2024flat ; jiang2024pressure ; chen2025charge ; yang_orbital-dependent_2024 ; NPzhang ; Hou_2023 ; PhysRevB.110.134520 ; wang_pressure-induced_2024 ; liu_electronic_2024 ; PhysRevB.110.L180501 ; PhysRevLett.134.076001 ; shao2024possiblehightemperaturesuperconductivitydriven . Subsequently, researchers observed SC in the trilayer nickelate La4Ni3O10 under pressure, and recently detected SC in ambient-pressure (AP) La3Ni2O7 thin films zhu_superconductivity_2024 ; zhang2025superconductivity ; ko_signatures_2025 ; zhou_ambient-pressure_2025 . These findings further underscore the potential of RP-series nickelates as high-$T_{c}$ superconductors PhysRevB.110.014503 ; PhysRevB.111.075140 ; PhysRevB.109.144511 ; li_signature_2024 ; hu2025electronicstructuresmultiorbitalmodels ; experimental . In RP nickelates with the general formula Rn+1NinO3n+1 (R = rare earth), each Ni forms NiO6 octahedron with corner-sharing O atoms. At AP, bilayer La3Ni2O7 and trilayer La4Ni3O10 manifest octahedral tilting and in-plane lattice anisotropy. Upon increasing pressure, both the octahedral tilting and the in-plane lattice constant disparity gradually diminish and eventually crystal structures arrive at higher symmetry ($I4/mmm$ space group) li_identification_2025 .
Furthermore, hybrid RP phase nickelates, formed by alternately stacking different RP phases along the c-axis, also attract widespread investigations, especially hybridization of single-layer La2NiO4, bilayer La3Ni2O7 and trilayer La4Ni3O10 zhang2024magneticcorrelationspairingtendencies ; chen_jpcm ; PhysRevLett.133.146002 ; PhysRevMaterials.8.053401 ; ouyang2025phasediagramskeyfactors . Until now, successful synthesis of 1313 phase (La3Ni2O7)  chen_jpcm ; PhysRevLett.133.146002 and 1212 phase (La5Ni3O11PhysRevMaterials.8.053401 has been achieved. A previous study reported a superconducting transition onset temperature up to 80 K in alternately stacking single-layer La2NiO4 and trilayer La4Ni3O10, namely 1313 phase PhysRevLett.133.146002 . However, it remains unclear whether the observed high-$T_{c}$ comes from the hybrid 1313 phase or $2$ bilayer phase.
Recently, pressure-induced SC with maximal $T_{c}=$ 64 K was observed in hybrid 1212 phase shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . Transport and magnetic torque measurements also revealed a density-wave (DW) transition at ambient and low pressures. In contrast to the previously reported pressure-induced SC in La3Ni2O7 and La4Ni3O10, the DW observed here is notably robust against pressure. Intriguingly, the superconducting phase emerges at around 11.7 GPa and exhibits a dome-like pressure-dependent behavior, in contrast to the previously reported right-triangle-like behavior observed in pressurized li_identification_2025 . Previous random-phase approximation (RPA) based study proposed that $d$-wave is the leading pairing symmetry that originates from the La2NiO4 single-layer in the material zhang2025electronicstructuremagneticpairing . However, recent studies using combined density functional theory (DFT) and dynamic-mean-field-theory (DMFT) provide another possibility that the La2NiO4 single-layer in La5Ni3O11 is nearly Mott-insulating which does not carry SC ouyang2025phasediagramskeyfactors ; labollita2025correlated , consistent with previous insight that the pure La2NiO4 is an antiferromagnetic (AFM) Mott insulator PhysRevB.75.012414 ; PhysRevB.80.144523 . Therefore, a more systematic investigation into the electronic and superconducting properties of the hybrid 1212 phase La5Ni3O11 is urgently needed.
In this paper, we adopt the first-principle DFT calculations to study the electronic structure of La5Ni3O11, followed by a RPA based study to clarify the superconducting mechanism in this material. Based on our DFT band structure, we construct a trilayer $(d_{z^{2}},d_{x^{2}-y^{2}})$- orbital tight-binding (TB) model, in which the hopping integrals between the bilayer subsystem and the single-layer one is very weak, reflecting that the two subsystems are almost decoupled, allowing for separate treatments of the two parts. Our RPA based analysis suggests that the superconducting pairing mainly occurs in the bilayer subsystem, which takes the s±-pattern similar with that in pressurized La3Ni2O7, while the sinlgle-layer subsystem mainly serves as a bridge connecting adjacent superconducting bilayers to establish phase coherence along the c-axis necessary for the bulk SC, through the interlayer Josephson coupling (IJC). As the IJC is extremely weak, it will experience prominent enhancement with the increase of the pressure, leading to enhancement of the $T_{c}$ in the low pressure regime. When the pressure is strong enough, the $T_{c}$ decreases with further enhancement of the pressure due to a reduction in the density of states (DOS) on the $\gamma$-pocket crucial for pairing. Therefore, our study provides a natural understanding of the dome-shaped pressure dependence of the $T_{c}$ observed in the experiment.
Results
Band Structure and TB model
Figure 1: (color online) Crystal structure. Side view of RP crystal structures of single-layer (La2NiO4), bilayer (La3Ni2O7) and SL-BL (La5Ni3O11). The blue, grey, and red balls represent lanthanum, nickel, and oxygen atoms.
La5Ni3O11 is a single-layer-bilayer (SL-BL) stacked nickelate, as shown in Fig. 1. La2NiO4 and La3Ni2O7 stack alternatively in RP phase, forming in $Cmmm$ space group at AP and transitioning to $P4/mmm$ space group under high pressure (HP) (shi2025superconductivityhybridruddlesdenpopperla5ni3o11, ). Unlike bulk La3Ni2O7, the out-of-plane Ni-O-Ni angle in La5Ni3O11 remains 180 from AP to HP, leading to the absence of octahedral distortion.
Band structure of SL-BL stacking La5Ni3O11 at 12 GPa is shown in Fig. 2(b), consistent with previous theoretical results ouyang2025phasediagramskeyfactors ; labollita2025correlated . As in most RP phase nickelates, bands around the Fermi level can be described by the Ni-$e_{g}$ sector. The Ni-$d_{z^{2}}$ band of La2NiO4 lies between the bonding and antibonding bands of La3Ni2O7. Compared to band structure at AP (see Fig.[S2]), the bonding state of La3Ni2O7 is metalized by pressure, leaving the $\gamma$ pocket at the Fermi surface (FS), as shown in Fig. 2(d). The emergence of $\gamma$ pocket implies that the SC in La5Ni3O11 may be attributed to La3Ni2O7, since the appearance of SC in La3Ni2O7 is also accompanied by the emergence of the $\gamma$ pocket.
Figure 2: (color online) Band structure of the DFT and six-orbital TB model for La5Ni3O11 at 12 GPa. (a) Schematic of La5Ni3O11 lattice of six-orbital TB model. The dashed line denotes the $(\frac{1}{2},\frac{1}{2})$ translation of RP stacking between bilayer and single-layer. (b) DFT band structure and projected DOS of La5Ni3O11. (c) The TB band structure corresponding to (a). (d) FS in the brillouin zone, with the five pockets labeled. The color scheme in the left half of (d) represents the relative contributions of the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbital, while the colors scheme on the right half of (d) indicates the relative contributions from Ni atoms in the single-layer and bilayer.
To further investigate the electronic properties in La5Ni3O11, we perform Wannier downfolding on the Ni-$d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals based on the DFT electronic structure. The obtained TB Hamiltonian in real space can be expressed as
$H_{{\rm TB}}=\sum_{\bm{r}_{i}\Delta\bm{r}\mu\nu\sigma}t_{\Delta\bm{r}\mu\nu}c^{\dagger}_{\bm{r}_{i}\mu\sigma}c_{(\bm{r}_{i}+\Delta\bm{r})\nu\sigma}.$ (1)
Here $\bm{r}_{i}$ represents the coordinates of site $i$, $\Delta\bm{r}_{x}(\Delta\bm{r}_{y})\in(-2,2)$ represents the hoppings up to the third-nearest neighbor. In this context, $\bm{r}_{i}$ refers to the unit cell index, while a finite $\Delta\bm{r}$ represents the interunit cell separation. The indices $\mu,\nu=1,\cdots,6$ containing the $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals of the upper and lower layers of the bilayer subsystem, as well as those of the single-layer subsystem. The SL-BL lattice and corresponding hoppings are depicted in Fig. 2(a), with the hopping integrals $t_{ij}^{\mu\nu}$ are provided in Tab. 1. Note that due to the RP stacking in La5Ni3O11, a $(\frac{1}{2},\frac{1}{2})$ translation exists between the bilayer and single-layer lattice. Fig. 2(c) displays the band structure for TB model, which is in good agreement with DFT results. The obtained FS for TB model reveals five pockets labeled as $\alpha$, $\alpha^{\prime}$, $\beta$, $\gamma$, $\gamma^{\prime}$, as shown in Fig. 2(d). The dominant component orbital of $\gamma$ and $\gamma^{\prime}$ pockets is $d_{z^{2}}$, while the $\alpha$, $\alpha^{\prime}$ and $\beta$ pockets are the mix of $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals. From the right half of Fig. 2(d), we can see that the $\alpha^{\prime}$ and $\gamma^{\prime}$ pocket are contributed from single-layer while the other three pockets from bilayer.
Table 1: Hopping parameters and site energies of TB model for La5Ni3O11 at 12 GPa in unit of eV. Here, $x$, $z$ and $xz$ denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intra-bilayer, inter-bilayer, and single-layer in-plane hoppings, respectively, while BS denotes hoppings between bilayer and single-layer units. The TB model includes hopping processes up to the third-nearest neighbor, i.e., [2,0]. $\epsilon$ denotes the on-site energy.
$t_{[1,0]}^{z,BI}$ $t_{[1,1]}^{z,BI}$ $t_{[1,0]}^{xz,BI}$ $t_{[1,0]}^{x,BI}$ $t_{[1,1]}^{x,BI}$ $t_{[0,0]}^{z,BO}$ $t_{[1,0]}^{z,BO}$
-0.1125 -0.0180 0.2373 -0.4708 0.0661 -0.6713 0.0202
$t_{[1,1]}^{z,BO}$ $t_{[0,0]}^{x,BO}$ $t_{[1,0]}^{x,BO}$ $t_{[1,1]}^{x,BO}$ $t_{[1,0]}^{xz,BO}$ $t_{[0,0]}^{z,BS}$ $t_{[1,0]}^{z,SI}$
0.0065 0.0132 -0.0013 0.0017 -0.0320 -0.0103 -0.0886
$t_{[1,1]}^{z,SI}$ $t_{[1,0]}^{xz,SI}$ $t_{[1,0]}^{x,SI}$ $t_{[1,1]}^{x,SI}$ $t_{[2,0]}^{z,BI}$ $t_{[2,0]}^{xz,BI}$ $t_{[2,0]}^{x,BI}$
-0.0108 0.1901 -0.4522 0.0752 -0.0192 0.0344 -0.0686
$t_{[2,0]}^{z,SI}$ $t_{[2,0]}^{xz,SI}$ $t_{[2,0]}^{x,SI}$ $\epsilon_{z}^{B}$ $\epsilon_{x}^{B}$ $\epsilon_{z}^{S}$ $\epsilon_{x}^{S}$
-0.0121 0.0256 -0.0590 0.3288 0.8648 0.1360 0.9313
Figure 3: (color online) DFT calculated band and FS characteristics of the bilayer subsystem and single-layer subsystem in La5Ni3O11 (SL-BL system) at 12 GPa. (a) Band structure of the SL-BL system (black lines) and the isolated bilayer (red lines). (b) FS of the bilayer subsystem, showing that the $\alpha$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\beta$ pocket contains contributions from both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, and the $\gamma$ pocket originates mainly from $d_{z^{2}}$ orbitals. (c) Band structure of the SL-BL system (black lines) and the isolated single-layer (red lines). (d) FS of the single-layer subsystem, showing that the $\alpha^{\prime}$ pocket is primarily composed of $d_{x^{2}-y^{2}}$ orbitals, the $\gamma^{\prime}$ pocket originates mainly from $d_{z^{2}}$ orbitals.
As shown in Tab. 1, the coupling between the single-layer and the bilayer is extremely weak, with the strongest hopping integral amplitude between the two parts to be $\left|t_{[0,0]}^{z,BS}\right|=0.0103$ eV. The sole effect of such a weak coupling is to equilibrate the chemical potentials of the two subsystems and thereby adjust the electron filling. Consequently, it is reasonable to treat the two subsystems separately.
The decoupling of the two subsystems at the TB Hamiltonian level is achieved by modifying the summation over $\mu$ and $\nu$ in Eq. 1. Specifically, restricting $\mu,\nu\in(1,\cdots,4)$ and electron filling $n\approx 3.1$ which got from SL-BL system yields the Hamiltonian of the bilayer subsystem, while restricting $\mu,\nu\in(5,6)$ and electron filling $n\approx 1.9$ yields that of the single-layer subsystem. A comparison between the band structure of the bilayer subsystem and that of La5Ni3O11 (SL-BL system) is shown in Fig. 3(a), where it can be seen that the bilayer subsystem reproduces the key features of the bilayer-derived bands in the full system. The three FS pockets of the bilayer, labeled $\alpha$, $\beta$, and $\gamma$, are shown in Fig. 3(b), and are clearly separated from the five pockets observed in Fig. 2(d). Fig. 3(c) shows that the band structure of the single-layer subsystem matches well with the single-layer-derived bands of the full SL-BL system. The two Fermi pockets of the single-layer subsystem, corresponding to the $\alpha^{\prime}$ and $\gamma^{\prime}$ pockets of the full system, are shown in Fig. 3(d).
Figure 4: (color online) (a) FS of the bilayer subsystem at 12 GPa. (b) Distribution of the largest eigenvalue of the spin susceptibility matrix $\chi^{s}(q)$ in the bilayer subsystem for $U=1$ eV and $J_{H}=U/6$. The susceptibility peaks at two inequivalent momenta, denoted as Q1 and Q2, respectively. As shown in (a), Q1 corresponds to a nesting vector between the $\beta$ and $\gamma$ pockets, while Q2 corresponds to a nesting vector between the $\alpha$ and $\beta$ pockets. (c) The largest pairing eigenvalue $\lambda$ of the various pairing symmetries as function of the interaction strength $U$ with fixed $J_{H}=U/6$. (d) The distributions of the leading $s^{\pm}$-wave pairing gap functions on the FS for $U=1.1$ eV.
The separation of the single-layer and bilayer subsystems within the TB framework points toward an underlying layer-selective character in this hybrid RP-phase nickelate. At ambient and low pressures, the incorporation of local dynamical correlations via the DMFT modifies the DFT band structure by removing the FS pockets contributed by the single‐layer subsystem, leaving only the three FS pockets contributed by the bilayer subsystem ouyang2025phasediagramskeyfactors ; labollita2025correlated . This observation indicates that the single-layer subsystem is close to a Mott‐insulating state, with possible AFM order present ouyang2025phasediagramskeyfactors , analogous to the pure La2NiO4 PhysRevB.75.012414 ; PhysRevB.80.144523 . At HP, the Mott gap becomes narrower, permitting a small density of mobile carriers that may suppress the AFM order ouyang2025phasediagramskeyfactors . Indeed, the long‐range magnetic order has not been observed experimentally under HP so far PhysRevMaterials.8.053401 . Regardless of whether residual magnetic order persists or not, the single-layer subsystem remains a “bad metal” and can hardly carry SC ouyang2025phasediagramskeyfactors . Consequently, SC must arise predominantly from the bilayer subsystem. Since DFT+DMFT indicates that electronic correlation in the bilayer subsystem is comparatively weaker, we will focus our subsequent RPA analysis on the bilayer subsystem to study its pairing instability.
RPA study of the SC
We adopt the following multi-orbital Hubbard interaction to investigate the SC drived by electron interactions in this SL-BL system,
$H_{int}$ $=U\sum_{i\tilde{\mu}}n_{i\tilde{\mu}\uparrow}n_{i\tilde{\mu}\downarrow}+V\sum_{i,\sigma,\sigma^{\prime}}n_{i1\sigma}n_{i2\sigma^{\prime}}$
$+J_{H}\sum_{i\sigma\sigma^{\prime}}\Big{[}c^{\dagger}_{i1\sigma}c^{\dagger}_{i2\sigma^{\prime}}c_{i1\sigma^{\prime}}c_{i2\sigma}+(c^{\dagger}_{i1\uparrow}c^{\dagger}_{i1\downarrow}c_{i2\downarrow}c_{i2\uparrow}+h.c.)\Big{]}.$ (2)
Here, $U$, $V$, and $J_{H}$ denote the intra-orbital, inter-orbital Hubbard repulsion, and the coupling of Hund (and the pair hopping) respectively, which satisfy the relation $U=V+2J_{H}$. $\tilde{\mu}\in(1,2)$ labels the two orbitals ($d_{z^{2}}$ and $d_{x^{2}-y^{2}}$) associated with each Ni atom. $i$ denotes the lattice sites belonging to the bilayer subsystem. We employ the multi-orbital RPA approach  takimoto2004strong ; yada2005origin ; kubo2007pairing ; graser2009near ; liu2013d+ ; zhang2022lifshitz ; kuroki101unconventional to treat this Hamiltonian. By renormalization, the spin susceptibility $\bm{\chi}^{(s)}$ and charge susceptibility $\bm{\chi}^{(c)}$ can be defined as Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11):
$\bm{\chi}^{(s)}(\bm{k},i\nu)=[I-\bm{\chi}^{(0)}(\bm{k},i\nu)U^{(s)}]^{-1}\bm{\chi}^{(0)}(\bm{k},i\nu),$
$\bm{\chi}^{(c)}(\bm{k},i\nu)=[I+\bm{\chi}^{(0)}(\bm{k},i\nu)U^{(c)}]^{-1}\bm{\chi}^{(0)}(\bm{k},i\nu).$ (3)
Where $\bm{\chi}^{(0)}$ is bare susceptibility for the non-interacting case, expressed as a tensor $\bm{\chi}^{(0)pq}_{st}$, with $p,q,s,t$ as orbital indices. Similarly, $\bm{\chi}^{(s)}$ can be expressed as $\bm{\chi}^{(s)pq}_{st}$. $U^{(s,c)}$ is the renormalized interaction strength, which is represented as a $4^{2}\times 4^{2}$ matrix in the bilayer subsystem. Note that there is a critical interaction strength $U_{c}^{(s/c)}$ for the spin/charge susceptibility. When $U\geq U_{c}^{(s,c)}$, the denominator matrix in Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11) will have zero eigenvalues for certain values of $\bm{k}$, causing the renormalized spin or charge susceptibility to diverge. This divergence indicates the onset of magnetic or charge order. Generally, repulsive Hubbard interactions suppress the charge susceptibility, but enhance the spin susceptibility  takimoto2004strong ; yada2005origin ; kubo2007pairing ; graser2009near ; liu2013d+ ; zhang2022lifshitz ; kuroki101unconventional . Therefore, $U_{c}^{(s)}<U_{c}^{(c)}$ and we denote $U_{c}^{(s)}$ as $U_{c}$ in the bilayer subsystem in the following. When fixing $J_{H}=U/6$, we find $U_{c}\approx 1.15$ eV.
Defining $\chi^{s}(\bm{q})$ as the maximum eigenvalue of $\bm{\chi}^{(s)}$ at each momentum, Fig. 4(b) shows its distribution in the brillouin zone for $U=1$ eV$<U_{c}$ of the bilayer subsystem. Notably, the distribution exhibits peaks at two unequivalent momenta, which we have labeled as $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$. These two momenta precisely correspond to the two FS nesting vectors of the bilayer subsystem, as illustrated in Fig. 4(a), and the vector $\mathbf{Q}_{1}=(\pi,0)$ is associated with the highest intensity of spin susceptibility, which is very similar to the case of the La3Ni2O7 superconductor under pressure WangQH2023 ; YangF2023 .
When $U<U_{c}$, the spin fluctuations can mediate SC in the bilayer subsystem, whose $T_{c}$ is related to the largest pairing eigenvalue $\lambda$ via $T_{c}\propto\omega_{D}e^{-1/\lambda}$ RPA5 , and the pairing symmetry is determined by the corresponding eigenvector (see $\bf Methods$). Fig. 4(c) illustrates the dependence of the largest pairing eigenvalue, denoted as $\lambda$, on the interaction strength $U$ for different potential pairing symmetries. The $D_{4h}$ point group of this bilayer structure allows for several possible pairing symmetries, in Fig. 4(c), we show the three leading pairing symmetries: $s$-wave, $d$-wave, and degenerate $(p_{x},p_{y})$-wave pairings. It is clear that the $s$-wave is the leading pairing symmetry and dominates other ones. The gap function of the obtained $s$-wave pairing is shown on the FS in Fig. 4(d), which displays the $s^{\pm}$ pattern. Consequently, the $\alpha$- and $\gamma$- pockets connected by the nesting vector $\mathbf{Q}_{1}$ are distributed with the strongest pairing amplitude, with their gap signs opposite. This pairing pattern is also similar to that in pressurized La3Ni2O7 luo2024high ; WangQH2023 ; YangF2023 .
Pressure Dependence of the SC
Figure 5: (color online) Pressure-dependence of the pairing eigen value $\lambda$ (black solid line) and the DOS on the $\gamma$-pocket (red solid line).
Table 2: Hopping parameters and site energies of TB model for La5Ni3O11 under various pressures. $x$, $z$ and $xz$ denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intrabilayer, interbilayer, and single-layer in-plane hoppings, respectively. Only hoppings greater than 0.1 eV are listed here for clarity. All parameters are in unit of eV.
Pressure (GPa) $\varepsilon_{z}^{B}$ $\varepsilon_{x}^{B}$ $\varepsilon_{z}^{S}$ $\varepsilon_{x}^{S}$ $t_{[0,0]}^{z,BO}$ $t_{[1,0]}^{x,BI}$ $t_{[1,0]}^{z,BI}$ $t_{[1,0]}^{xz,BI}$ $t_{[1,0]}^{x,SI}$ $t_{[1,0]}^{xz,SI}$
12.46 0.3343 0.9145 -0.1624 1.0219 -0.6727 -0.4773 -0.1134 0.2412 -0.4595 0.1872
15.82 0.3352 0.9210 -0.1634 1.0241 -0.6948 -0.4869 -0.1222 0.2452 -0.4691 0.1921
19.40 0.3369 0.9219 -0.1852 1.0270 -0.7239 -0.5009 -0.1224 0.2511 -0.4783 0.1955
23.28 0.3370 0.9295 -0.1877 1.0290 -0.7492 -0.5105 -0.1279 0.2557 -0.4894 0.1999
27.39 0.3375 0.9311 -0.2067 1.0291 -0.7592 -0.5230 -0.1314 0.2610 -0.5003 0.2031
31.82 0.3379 0.9409 -0.2099 1.0315 -0.7817 -0.5339 -0.1358 0.2659 -0.5103 0.2080
36.51 0.3386 0.9422 -0.2213 1.0357 -0.8005 -0.5464 -0.1422 0.2710 -0.5225 0.2117
41.54 0.3402 0.9506 -0.2337 1.0350 -0.8124 -0.5574 -0.1458 0.2754 -0.5335 0.2156
To investigate the pressure dependence of SC in the system, we first performed DFT calculations of the electronic band structure under pressures ranging from approximately 12 to 40 GPa. The corresponding key hopping parameters at each pressure are listed in Tab. 2. Based on these TB results, we employed the RPA approach to calculate the pairing eigenvalue $\lambda$ as a function of pressure $P$. As shown in Fig. 5, the resulting $\lambda$ exhibits a downward trend with enhanced $P$, which originates from reduced DOS on the $\gamma$ pocket.
The decreasing $\lambda\sim P$ relation caused by the reduced DOS shown in Fig. 5 parallels that in the pure La3Ni2O7 PhysRevLett.134.076001 ; huomodulation ; zhang2024structural , which closely matches the experimentally observed $T_{c}\sim P$ relation in that material li_identification_2025 . However, it seems at a glance that such a decreasing $\lambda\sim P$ relation conflicts with the $T_{c}\sim P$ relation in La5Ni3O11, which shows a pronounced increase up to around 20 GPa, followed by a gradual decline afterward shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . The discrepancy between the $\lambda\sim P$ and the $T_{c}\sim P$ relations in La5Ni3O11 originates from that its alternating SL-BL structure brings into extremely weak IJC that is crucial for the establishment of the bulk SC, which can be prominently enhanced by the enhancement of $P$. In the following, we clarify this viewpoint.
Figure 6: (color online) Influence of the IJC on the pressure-dependence of the $T_{c}$. (a) Schematic of the intrinsic S–N–S Josephson Junction structure of La5Ni3O11, where the superconducting (S) layers correspond to the La3Ni2O7 bilayer and the normal (N) layers correspond to the La2NiO4 single-layer. (b) Schematic of the interlayer distance $d_{\perp}$ dependence of the hopping integral $t_{z}^{BS}$ between the bilayer subsystem and the single-layer subsystem. (c) $T_{c}/\rho_{0}$ as a function of the anisotropy parameter $\eta$ that reflects the strength of the IJC. The red dashed line marks the $\eta\approx\mathrm{1}0^{-8}$ associated with the La5Ni3O11, while the blue dashed line marks the $\eta\approx\mathrm{1}0^{-4}$ associated with the La3Ni2O7. (d) Pressure-dependence of the superconducting $T_{c}$ (in unit of $\omega_{D}$) in La5Ni3O11.
As shown in Fig. 6(a), the La5Ni3O11 hosts alternating bilayer and single-layer NiO2 subsystems, where the former supports SC and the latter is a “bad metal”. This results in an intrinsic superconductor–normal metal–superconductor (S–N–S) Josephson Junction structure. In this structure, the establishment of bulk SC requires not only intra-bilayer pairing but also inter-bilayer phase coherence. Physically, the inter-bilayer phase coherence is established via combined tunneling of a Cooper pair from a superconducting bilayer to its adjacent one, i.e. the IJC, across the intervening normal-metallic single-layer. This is a second-order process of the inter-bilayer single-particle tunneling, and hence a fourth-order process of the weak bilayer-single-layer tunneling for a single particle. As such an IJC is extremely weak, it will experience a prominent enhancement with the increase of the pressure. In the following, we provide a more quantitative clarification for this point.
In principle, the hopping integral $t_{z}^{BS}$ between the bilayer and the single-layer subsystems can be obtained through fitting the TB model to the DFT band structure. However, due to the limited accuracy achieved in the fitting, the obtained tiny $t_{z}^{BS}$ bears a considerably large relative error bar, and the accurate pressure dependence of $t_{z}^{BS}$ can be hardly acquired. To overcome this difficulty, we adopt the following semi-phenomenological analysis for the problem. As shown in Fig. 6(b), the $t_{z}^{BS}$ depends on the overlap between the $d_{z^{2}}$ orbital wavefunctions of the bilayer and the single-layer subsystems, which decays exponentially with the displayed interlayer distance $d_{\perp}$, i.e. $t_{z}^{BS}\sim e^{-d_{\perp}/\xi}$, where $\xi$ characterizes the decaying length. Consequently, the effective hopping integral between adjacent superconducting bilayers follows $t_{z}^{BB}\sim(t_{z}^{BS})^{2}$, and the IJC follows $J_{\perp}^{BB}\propto(t_{z}^{BB})^{2}\sim(t_{z}^{BS})^{4}\sim e^{-4d_{\perp}/\xi}$.
In a quasi-2D layered system such as La5Ni3O11 and La3Ni2O7, the bulk $T_{c}$ is related to the anisotropy parameter $\eta$ through the following relation in the limit of $\eta\to 0$ kopec2000superconducting ,
$T_{c}\approx\rho_{0}\frac{\pi}{\mathrm{ln}(32/\eta)}.$ (4)
Here $\rho_{0}$ indicates the intralayer phase stiffness reflecting the pairing within the bilayer subsystem, which is approximated as the pairing temperature obtained from the RPA calculation, i.e. $\omega_{D}e^{-1/\lambda}$. Here the “Debye frequency” $\omega_{D}$ represents for the typical energy scale of spin fluctuations. The anisotropy parameter $\eta$ is the ratio between the IJC strength $J_{\perp}^{BB}$ and $\rho_{0}$. Since $J_{\perp}^{BB}\propto(t_{z}^{BS})^{4}$, and $t_{z}^{BS}\sim 10^{-2}$, we have $\eta\sim 10^{-8}$ for La5Ni3O11, which is extremely weak. In contrast, for La3Ni2O7, since there exists no intervening normal-metallic single-layer to go across, we have $\eta\sim 10^{-4}$. As shown in Fig. 6(c), for La5Ni3O11, the anisotropy parameter $\eta$ locates within the regime that a slight enhancement of $\eta$ will prominently enhance the $T_{c}$; while for La3Ni2O7, $\eta$ has only a minor effect on $T_{c}$.
Under applied pressure, the strain is proportional to the pressure, implying that $\Delta d_{\perp}\propto-P$, i.e., $d_{\perp}(P)=d_{\perp}^{0}-\kappa P$, where $d_{\perp}^{0}$ is the interlayer distance at zero pressure and $\kappa$ is a proportionality constant. Therefore,
$\eta=e^{-4(d_{\perp}^{0}-\kappa P)/\xi}\equiv\eta_{0}e^{\alpha P}.$ (5)
Here, $\eta_{0}$ is the anisotropy parameter at AP. Ultimately,
$T_{c}\approx T_{c}^{\mathrm{RPA}}\frac{\pi}{\mathrm{ln}(\frac{32}{\eta})}\approx\omega_{D}e^{-1/\lambda}\frac{\pi}{\mathrm{ln}(\frac{32}{\eta_{0}e^{\alpha P}})}.$ (6)
For a reasonable set of parameters, e.g. $\eta_{0}\approx 10^{-8}$ and $\alpha\approx 0.3$ GPa-1, the resulting pressure dependence of $T_{c}$ shown in Fig. 6(d) indeed exhibits a dome-shaped behavior similar to that observed in experiment shi2025superconductivityhybridruddlesdenpopperla5ni3o11 .
Discussion
In this work, we investigate the electronic structure and superconducting mechanism of the hybrid RP nickelate La5Ni3O11, composed of alternating single-layer and bilayer subsystems. Based on DFT and RPA calculations, we find that SC predominantly emerges in the bilayer subsystem with an $s^{\pm}$-wave pairing symmetry, closely resembling that in pressurized La3Ni2O7. We further show that the pressure dependence of the superconducting $T_{c}$ in La5Ni3O11 is governed not only by the intralayer pairing but also by the IJC, which leads to a dome-like $T_{c}(P)$ behavior consistent with experimental observations.
Methods
DFT method
DFT calculations were performed by Vienna ab initio simulation package (VASP) (VASP1, ; VASP2, ), in which the projector augmented wave (PAW) (PAW1, ; PAW2, ) method with a 600 eV plane-wave cutoff is applied. The generalized gradient approximation (GGA) of Perdew-Burke-Ernzerhof (PBE) form exchange correlation potential is adopted  (PhysRevLett.77.3865, ). The convergence criterion of force was set to be 0.001 eV/Å and total energy convergence criterion was set to be $10^{-7}$ eV. A $\Gamma$-centered $19\times 19\times 5$ Monkhorst Pack k-mesh grid is used for primitive cell of $P4/mmm$ phase and a $\Gamma$-centered $14\times 14\times 5$ Monkhorst Pack k-mesh grid for primitive cell of $Cmmm$ phase. In structural relaxations, the lattice constants of $P4/mmmm$ phase we use are $a=3.77$  and $c=16.20$. DFT calculations reveal that the corresponding pressure is 12 GPa, where La5Ni3O11 enters superconducting state  shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . An effective Hubbard U parameter $U=3.5$ eV was employed to account for the correlation effects of 3d electrons in Ni atoms (yang_orbital-dependent_2024, ; PhysRevB.57.1505, ). To obtain the projected TB models, we further performed Wannier downfolding as implemented in WANNIER90 (w90, ) package, in which the good convergences were reached.
RPA method
In the standard multi-orbital RPA approach, the bare susceptibility is defined as
$\chi^{(0)pq}_{st}(\bm{k},\tau)\equiv$ $\frac{1}{N}\sum_{\bm{k}_{1}\bm{k}_{2}}\left\langle T_{\tau}c_{p}^{\dagger}(\bm{k}_{1},\tau)c_{q}(\bm{k}_{1}+\bm{k},\tau)\right.$
×$\left.\times c_{s}^{\dagger}(\bm{k}_{2}+\bm{k},0)c_{t}(\bm{k}_{2},0)\right\rangle_{0},$ (7)
where $\langle\cdots\rangle_{0}$ represents the expectation value in the free-electron state, and $p/q/s/t$ are the effective orbital indices, which label combined layer, sublattice and physical orbital indices. Transforming the above defining formula to the momentum-frequency space, we obtain the explicit formula of bare susceptibility as
$\chi^{(0)pq}_{st}$ $(\bm{k},i\omega_{n})=\frac{1}{N}\sum_{\bm{k}_{1}\alpha\beta}\xi^{*}_{\alpha p}(\bm{k}_{1})\xi_{\beta q}(\bm{k}_{1}+\bm{k})\xi^{*}_{\beta s}(\bm{k}_{1}+\bm{k})$
$\xi_{\alpha t}(\bm{k}_{1})\frac{f(\varepsilon^{\beta}_{\bm{k}_{1}+\bm{k}}-\mu_{c})-f(\varepsilon^{\alpha}_{\bm{k}_{1}}-\mu_{c})}{i\omega_{n}+\varepsilon^{\alpha}_{\bm{k}_{1}}-\varepsilon^{\beta}_{\bm{k}_{1}+\bm{k}}},$ (8)
where $\alpha,\beta$ represent band indices, $\varepsilon$ and $\xi$ are the eigen-value, and eigen-state of the free particle Hamiltonian, $\mu_{c}$ is the chemical potential, and $f(\varepsilon_{\bm{k}})=1/(1+e^{\beta\varepsilon_{\bm{k}}})$ is the Fermi-Dirac function.
In the RPA level, the renormalized spin/charge susceptibilities for the system are
$\chi^{(s)}(\mathbf{q},i\nu)=[I-\chi^{(0)}(\mathbf{q},i\nu)U^{(s)}]^{-1}\chi^{(0)}(\mathbf{q},i\nu),$
$\chi^{(c)}(\mathbf{q},i\nu)=[I+\chi^{(0)}(\mathbf{q},i\nu)U^{(c)}]^{-1}\chi^{(0)}(\mathbf{q},i\nu).$ (9)
Where $\chi^{(s,c)}(\mathbf{q},i\nu)$, $\chi^{(0)}(\mathbf{q},i\nu)$, and $U^{(s,c)}$ are operated as $l^{2}$$\times$ $l^{2}$ matrices ($l$ represents the number of orbits and the upper or lower two indices are viewed as one number) with elements of the matrix $U^{(s/c)}$ to be
$U^{(s)pq}_{st}=\left\{\begin{aligned} U,~{}~{}~{}p=q=s=t\\ J_{H},~{}~{}~{}p=q\neq s=t\\ J_{H},~{}~{}~{}p=s\neq q=t\\ V,~{}~{}~{}p=t\neq s=q\\ \end{aligned}\right.$
$U^{(c)pq}_{st}=\left\{\begin{aligned} U,~{}~{}~{}p=q=s=t\\ 2V-J_{H},~{}~{}~{}p=q\neq s=t\\ J_{H},~{}~{}~{}p=s\neq q=t\\ 2J_{H}-V,~{}~{}~{}p=t\neq s=q\\ \end{aligned}\right.$
Since we only consider the on-site interaction, the elements of the matrix $U^{(s/c)}$ are non-zero only if $p,q,s,t$ are the same layer indices.
Note that there is a critical interaction strength $U_{c}^{(s,c)}$ for spin and charge respectively. When $U>U_{c}^{(s,c)}$, the denominator matrix in Eq. Pairing mechanism and superconductivity in pressurized La5Ni3O11 will have zero eigenvalues for some $\mathrm{q}$ and the renormalized spin or charge susceptibility diverges there, which invalidates the RPA treatment RPA1 ; RPA2 ; RPA3 ; RPA4 ; RPA5 ; RPA6 ; RPA7 ; RPA8 ; PhysRevB.110.L180501 . This divergence of spin susceptibility for $U>U_{c}^{(s)}$ implies magnetic order, while that of the charge susceptibility for $U>U_{c}^{(c)}$ implies charge order.
When $U<U_{c}$, a Cooper pair $c_{t}(\mathbf{q})c_{s}(-\mathbf{q})$ could be scattered to $c_{p}^{\dagger}(\mathbf{k})c_{q}^{\dagger}(-\mathbf{k})$ by exchanging charge or spin fluctuations. Considering only intra-band pairings, we obtain the following effective pairing interaction:
$V_{eff}=\frac{1}{N}\sum_{\alpha\beta\mathbf{k}\mathbf{q}}V^{\alpha\beta}(\mathbf{k},\mathbf{q})c_{\alpha}^{\dagger}(\mathbf{k})c_{\alpha}^{\dagger}(-\mathbf{k})c_{\beta}(-\mathbf{q})c_{\beta}(\mathbf{q})$ (10)
Here $\alpha$/$\beta$ is band indice. And effective pairing interaction vertex $V^{\alpha\beta}(\mathbf{k},\mathbf{q})$ has the form:
$V^{\alpha\beta}(\mathbf{k},\mathbf{q})=\sum_{pqst}\Gamma_{pq}^{st}(\mathbf{k},\mathbf{q})\xi^{\alpha,*}_{p}(\mathbf{k})\xi_{q}^{\alpha,*}(-\mathbf{k})\xi_{s}^{\beta}(-\mathbf{q})\xi_{t}^{\beta}(\mathbf{q}).$ (11)
Here $\xi^{\alpha}\left(\mathbf{k}\right)$ are the eigenvector corresponding to $\alpha$-th eigenvalue (relative to the chemical potential $\mu_{c}$) of the TB Hamiltonian matrices. Within the mean-field approximation, one can derive a self-consistent equation for the pairing gap, which when linearized near $T_{c}$ RPA1 ; RPA2 ; RPA3 ; RPA4 ; RPA5 ; RPA6 ; RPA7 ; RPA8 ; PhysRevB.110.L180501 yields the linearized gap equation (12):
$\Delta_{\alpha}(\mathbf{k})=-\sum_{\beta\mathbf{q}}V^{\alpha\beta}(\mathbf{k},\mathbf{q})\times\frac{\tanh(\frac{\beta_{c}}{2}\left|\tilde{\varepsilon}_{\beta}(\mathbf{q})\right|)}{\left|\tilde{\varepsilon}_{\beta}(\mathbf{q})\right|}\times\Delta_{\beta}(\mathbf{q}).$ (12)
Here $\beta_{c}$=$\frac{1}{\mathrm{k_{B}}T_{c}}$ represents the critical temperature of SC. Choosing a thin energy shell near the Fermi level, Eq. (12) becomes the eigenvalue problem of the effective interaction matrix $V^{\alpha\beta}$, which determines the $T_{c}$ and the leading pairing symmetry of the system. To be specific, $\Delta_{\alpha}(\mathbf{k})$ represents the relative gap function on the $\alpha$-th FS patches near $T_{c}$, and eigenvalue $\lambda$ is related to $T_{c}$ through $\lambda^{-1}=\mathrm{ln}(1.13\frac{\hbar\omega_{D}}{k_{\mathrm{B}}}T_{c})$. The leading pairing symmetry is determined by the largest eigenvalue $\lambda$ of Eq. (13):
[…]
Data availability
Relevant data supporting the key findings of this study are available within the article and the Supplementary Information file. All raw data generated during the current study are available from the corresponding authors upon reasonable request.
Code availability
The code that supports the plots within this paper is available from the corresponding author upon reasonable request.
Acknowledgements
We are grateful for the discussion with Chen Lu. This work is supported by National Natural Science Foundation of China (Grants No. 12494591, No. 92165204, No. 12234016, No. 12074031), National Key Research and Development Program of China (2022YFA1402802), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010). Ming Zhang is supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No. ZCLQN25A0402.
Author contribution
Dao-Xin Yao and Fan Yang conceived and designed the project. Ming Zhang and Cui-Qun Chen performed the numerical calculations for DFT, TB model, and RPA calculations. All authors contributed to the discussion of the results and wrote the paper.
Competing interests
The authors declare no competing interests.
Addition information
Supplementary information The online version contains supplementary information available at